An athlete threw a basketball a distance of 27.5m to score and win the game. If the shot was made at a 50.0 degree angle above the horizontal, what was the initial speed of the ball? (16.6m/s)
1
Expert's answer
2019-07-09T11:23:46-0400

We can find the initial speed of the ball from the projectile range equation:


"R = \dfrac{v_0^2sin2\theta}{g},"

here, "R = 27.5 m" is the range of the projectile, "v_0" is the initial speed of the ball, "\theta = 50^{\circ}" is the launch angle and "g = 9.8 m\/s^2" is the acceleration due to gravity.

Then, from this formula we can calculate the initial velocity of the ball:


"v = \sqrt{\dfrac{Rg}{sin2\theta}} = \sqrt{\dfrac{27.5 m \cdot 9.8\dfrac{m}{s^2}}{sin2 \cdot 50^{\circ}}} = 16.6 \dfrac{m}{s}."

Answer:

"v_0 = 16.6 \dfrac{m}{s}."